OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@zepdrix

OpenStudy (haseeb96):

-(3)^2 -3(2)^3 +(-3)^2 + 2(2)^3 -(-3)^2 now just solve it -9-3(8) + 9 + 2 (8) - 9 -24 +16 - 9 -17 is the correct answer x^2 = x multiply with x or 2^2 = 2 into 2 =4

OpenStudy (anonymous):

Can you help with one more?

OpenStudy (haseeb96):

yeah

OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2.

OpenStudy (haseeb96):

just put all these terms in it here it is not OP told us to solve the question i have already tell u how to solve it

OpenStudy (anonymous):

No i keep getting wierd answers

OpenStudy (anonymous):

I am not doing something right

OpenStudy (anonymous):

@Haseeb96 please help

OpenStudy (anonymous):

@ganeshie8 can you help

OpenStudy (haseeb96):

just first solve like me put the values of x,y and z in the given equation

OpenStudy (anonymous):

ok can you do it also with me

OpenStudy (haseeb96):

yes but first put these values in the equation

OpenStudy (anonymous):

ok id did

OpenStudy (anonymous):

\[2-4 ^{2}-2x 2^{2}+3^{2}+2^{2}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

2-16-8+9+4

OpenStudy (haseeb96):

2(-4)^2 -2 (2)^4 + (3)^2 -(-4)^2 + (2)^4 this is correct

OpenStudy (anonymous):

is what I last said right

OpenStudy (haseeb96):

now solve it

OpenStudy (haseeb96):

no

OpenStudy (anonymous):

is it 9

OpenStudy (haseeb96):

2(16) -2 (16) +9 -16 +16 the correct answer is 9

OpenStudy (anonymous):

Yay ill tell you if i get it right

OpenStudy (anonymous):

yes

OpenStudy (haseeb96):

okay but good job boy